### v4

# SpeechRole: A Large-Scale Dataset and Benchmark for Evaluating Speech Role-Playing Agents

Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
[…]
Figure 1: Overview of the SpeechRole framework. It covers role extraction with distinctive timbre and prosody, speech dialogue data construction, cascaded and end-to-end speech generation paradigms, and the SpeechRole-Eval benchmark for multidimensional evaluation of interaction abilities, speech quality, and role-playing fidelity.
1 Introduction
Recently, Role-Playing Agents (RPAs) have emerged as a promising paradigm for achieving personalized interactions and emotional resonance [DBLP:journals/tmlr/Chen00YZSXLYZCL24]. Simulated characters with distinct traits make interactions more engaging in applications such as digital assistants [DBLP:journals/corr/abs-2407-05305], educational tools [DBLP:journals/corr/abs-2409-03512], and interactive storytelling [DBLP:conf/acl/WangPQLZWGGN00024]. However, current research predominantly focuses on the text modality [DBLP:journals/corr/abs-2308-09597, DBLP:conf/emnlp/ShaoLDQ23, DBLP:conf/acl/WangPQLZWGGN00024, DBLP:conf/emnlp/SadeqXKLGM24], overlooking the crucial role of speech in authentic interactions.
Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [DBLP:conf/iui/ChangSZJ0K23]. Although preliminary studies have explored role-playing incorporating visual or speech modalities [DBLP:conf/iclr/DaiH0J0025, DBLP:journals/corr/abs-2505-20277], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs). SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions. In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
To address this gap, we propose SpeechRole, a unified framework for the development and evaluation of SRPAs (see Figure 1). We first introduce SpeechRole-Data, a large-scale dataset comprising 112k speech-based dialogues covering 98 diverse roles. Each role features distinctive timbre and prosodic characteristics, enabling expressive and role-consistent speech interactions.
Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses agents across three dimensions: interaction competence, speech expressiveness, and role consistency. This benchmark supports both cascaded and end-to-end architectures, facilitating comprehensive and comparable analyses across different paradigms.
Our experiments reveal complementary strengths and weaknesses between SRPA paradigms. Cascaded systems offer strong modular control and stable role identity across interactions, but may suffer from cross-module mismatches or limited expressiveness. In contrast, end-to-end models produce more coherent and natural speech, yet often struggle with speaker consistency and preserving character traits over extended dialogues. These trade-offs highlight the unique challenges in building effective SRPAs that balance interaction ability, speech expressiveness, and role-playing fidelity.
Overall, our main contributions are as follows:
  1. 1.
    We construct SpeechRole-Data, a large-scale multilingual corpus covering 98 roles and 112k speech-based dialogues with rich prosodic and timbral variation.
  2. 2.
    We develop SpeechRole-Eval, a comprehensive benchmark designed to evaluate SRPA capabilities across key dimensions including interaction ability, speech expressiveness, and role-playing fidelity.
  3. 3.
    We systematically compare cascaded and end-to-end paradigms, revealing their strengths, limitations, and offering insights for future SRPA development.
2 Related Work
Role-Playing Agents Across Modalities.
Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality. ChatHaruhi [DBLP:journals/corr/abs-2308-09597] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], Character-LLM [DBLP:conf/emnlp/ShaoLDQ23], and RoleFact [DBLP:conf/emnlp/SadeqXKLGM24] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations. Beyond text, MMRole [DBLP:conf/iclr/DaiH0J0025] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the value of speech for real-time, voice-aware character interaction.
To assess role-playing capabilities, benchmarks such as RoleEval [DBLP:journals/corr/abs-2312-16132], RoleInteract [DBLP:journals/corr/abs-2403-13679], SocialBench [DBLP:conf/acl/ChenCYXXSQLZH24], and CharacterEval [DBLP:conf/acl/TuFTSSGY24] evaluated behavioral consistency and persona fidelity, while INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24] and ECHO [DBLP:journals/corr/abs-2404-13957] introduced human-aligned, psychometric evaluation paradigms. These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity. However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
Interactive Speech Agents.
In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24]. Modern TTS models like F5TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], E2-TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], Spark-TTS [DBLP:journals/corr/abs-2503-01710], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025], and XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24] have significantly improved output quality, expressiveness, and controllability. However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
End-to-end systems aim to unify speech perception and generation within a single model. Early efforts like SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [DBLP:journals/corr/abs-2306-12925] utilized vocoder-based representations for continuous speech modeling. Recent advances, including Moshi [DBLP:journals/corr/abs-2410-00037], LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025], and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
Figure 2: Examples of the three major task categories in SpeechRole.
3 SpeechRole-Data
We introduce SpeechRole-Data, a large-scale dataset for speech-based role-play, featuring 98 roles and over 112k dialogues with reference audio and rich character profiles. To ensure personality and voice consistency, we follow a structured pipeline spanning role selection, dialogue generation, and audio processing. Details are in the appendix.
3.1 Text Data Construction
Role Selection.
We curate 98 diverse roles, including 18 from ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and 80 from RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024]. The selection balances language, gender, and personality, focusing on characters with distinctive vocal traits to enrich speech modeling.
Role Metadata Extraction.
For each character, we extract structured metadata to build a comprehensive profile, including (1) Role Profile (temperament, preferences), (2) Background (social identity, relationships), and (3) Character Lines (dialogues, monologues). This metadata forms the basis for generating personality-consistent dialogues.
Dialogue Generation.
Using the extracted metadata as a guide, we employ gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate conversations. For each of the 98 roles, we produce approximately 800 single-turn and 800 multi-turn dialogues, creating a substantial initial pool of text data.
Dialogue Filtering and Deduplication.
The generated data undergoes a multi-stage filtering process. First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese. To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
Splits Characters Samples (single-turn/multi-turn) Speech Hours (user/character)
Train 78 89,461 (43,082/46,379) 875.25 (204.45/670.80)
Dev 20 21,993 (10,627/11,366) 203.60 (49.84/153.76)
Test 98 392 (196/196) 4.20 (0.94/3.26)
Table 1: Statistics of SpeechRole-Data across train, dev, and test splits.
3.2 Role Voice Collection
Voice Collection and Preprocessing.
To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works. All audio is anonymized and fragmented into brief utterances to comply with fair use. Raw streams are extracted using ffmpeg [ffmpeg] and converted to mono 16 kHz WAV format.
Voice Cleaning and Annotation.
We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, including source separation, speaker diarization, and voice activity detection. This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23] and scored with DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22].
Figure 3: Speech duration distribution of users and roles.
Speaker Identification.
The diarization process outputs speaker indices without explicit identity information. To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to analyze transcribed text. The models’ integrated results are then manually verified.
Reference Audio Ranking and Selection.
To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity. We use the CAM++ model [DBLP:conf/interspeech/WangZCC023] to extract a speaker embedding for each clip. The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
3.3 Dataset Statistics
Dataset Overview.
SpeechRole-Data comprises 98 distinct roles and 112k dialogue samples, including both single-turn and multi-turn conversations (see Table 1). Each role is accompanied by a detailed role profile and a reference audio clip that demonstrates its unique speaking style, such as timbre, rhythm, and prosody. Each dialogue sample contains both text and speech, along with contextual background information relevant to the role, thereby supporting realistic, evidence-based role-playing. During the evaluation phase, the test set consists of two single-turn and two multi-turn dialogues sampled for each role. The remaining data are divided into a training set (78 roles) and a validation set (20 roles). Detailed examples are provided in Figure 2.
Speech Duration.
We analyze the speech duration of users and roles. As shown in Figure 3, the duration of role utterances is generally longer, typically ranging from 10 to 20 seconds. This is mainly due to the narrative or expressive nature of their content. In contrast, user utterances are concentrated within 3 to 6 seconds, reflecting their concise and goal-oriented characteristics.
4 SpeechRole-Eval
To systematically assess SRPAs, we introduce SpeechRole-Eval, a comprehensive benchmark that evaluates model performance across three key dimensions: fundamental interaction ability, speech expressiveness, and role-playing fidelity. In contrast to prior work that primarily focuses on text, SpeechRole-Eval places special emphasis on the speech modality as a critical component of evaluation. The scoring process is fully automated, powered by a strong multimodal model, and has been validated against human judgments to ensure reliability. This section describes the task setup, evaluation metrics, and the overall evaluation pipeline.
4.1 Task Types
To assess the cognitive, affective, and social competencies of SRPAs, user prompts are categorized into three major task categories based on communicative intent and interactional focus: (1) Internal Reasoning, which concerns the character’s rich subjective world, including personal perspectives, emotional responses, and underlying motivations; (2) Experiential Narration, which emphasizes the character’s ability to vividly recall past events, describe concrete actions, and provide relevant background or historical information; and (3) Social Communication, which highlights interpersonal understanding achieved through nuanced evaluative judgments, the provision of advice, and relational reasoning in context. These categories reflect the core dimensions of role-playing dialogue and enable a comprehensive, structured assessment of role-playing abilities (see Figure 2).
4.2 Evaluation Metrics
In contrast to traditional text-based RPAs, SRPAs must not only accurately emulate specific characters but also consistently deliver coherent and expressive spoken responses fully aligned with both the dialogue context and the character’s unique persona. To comprehensively evaluate SRPAs, we propose a robust three-dimensional evaluation framework that assesses fundamental interaction ability, speech quality and expressiveness, and role-playing fidelity.
Fundamental interaction abilities comprehensively evaluate the agent’s capacity to maintain coherent, natural, and instruction-aligned interactions fully consistent with the character, which are assessed by the following key metrics:
  • Instruction Adherence (IA): Do the spoken responses strictly follow the task instruction, remaining fully in character without any out-of-role explanations or assistant-like meta-comments?
  • Speech Fluency (SF): Are the responses delivered fluently, with smooth articulation, appropriate pacing, and minimal disfluencies such as stuttering or unnatural pauses?
  • Conversational Coherence (CC): Do the responses maintain logical consistency within the dialogue, aligning with previous content without contradictions or abrupt topic shifts?
Speech quality and expressiveness measure the naturalness and prosodic expressiveness of the generated speech, which are assessed by the following metrics:
  • Speech Naturalness (SN): Do the responses sound natural, human-like, and free from noticeable artifacts or robotic effects typically associated with synthetic speech?
  • Prosodic Consistency (PC): Does the prosody, including pitch, stress, and intonation, align with the character’s intended speaking style and remain consistent across the discourse?
  • Emotion Appropriateness (EA): Are emotional cues in the speech (e.g., anger, joy, sadness) well-aligned with the dialogue context and the character’s emotional state?
Role-playing fidelity assesses the extent to which the agent embodies the assigned character’s persona, which is assessed by the following metrics:
  • Personality Consistency (PeC): Do the responses consistently reflect the character’s personality traits, such as optimism, sarcasm, or authority?
  • Knowledge Consistency (KC): Are the responses grounded in the character’s established background, knowledge, and relationships, without fabricating out-of-character facts?
4.3 Evaluation Procedure
LLM-Based Scoring Procedure.
There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [DBLP:conf/acl/WangLCCZLCKLLS24]. Inspired by DBLP:conf/iclr/DaiH0J0025, two strategies are adopted to enhance the reliability of LLM-based scoring.
First, the model is required to generate a brief rationale for its score prior to providing the original rating. This process encourages more deliberate and interpretable evaluations by the model. Second, rather than evaluating individual speech samples in isolation, a paired comparison setting is designed. For each evaluation instance, the model receives a test speech sample and a high-quality reference answer simultaneously, and assigns scores to both.
The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies. This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments. We use gemini-2.5-pro-preview-06-05 [DBLP:journals/corr/abs-2312-11805] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
Human Correlation Verification.
To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses. Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric. The final human score for each metric is calculated as the average of all annotators’ judgments. The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation. The human judgment setting follows the evaluation protocol introduced by DBLP:conf/iclr/DaiH0J0025, which uses pairwise preference comparisons to ensure stable and interpretable annotations. Details are in the appendix.
Models Evaluation Metrics Overall
IA SF $C$ SN PC EA PeC KC
English Evaluation Results
Alibaba Cloud API 0.9012 0.6678 1.0320 0.6973 0.6651 0.8087 0.8594 0.9171 0.8186
XTTS 0.9617 0.5493 0.9872 0.6578 0.7568 1.0133 1.0357 0.9531 0.8644
CosyVoice2 1.0102 0.7869 1.0346 0.8682 0.9857 1.0781 1.0476 0.9744 0.9732
E2 TTS 1.0531 0.7938 1.1621 0.8859 1.0718 1.1806 1.1760 1.0104 1.0417
F5-TTS 1.1245 0.8057 1.1132 0.8717 0.9773 1.2686 1.0952 1.0614 1.0397
FireRedTTS 0.9587 0.7266 1.0297 0.8135 0.8760 1.0557 1.0415 0.9899 0.9365
Spark-TTS 1.0384 0.7583 1.0641 0.8282 1.0554 1.1427 1.0741 1.0144 0.9970
GPT-SoVITS 0.8566 0.4783 0.9293 0.5576 0.5606 0.8253 0.8899 0.8770 0.7468
MaskGCT 0.9277 0.7635 0.9902 0.7589 0.9449 1.0382 0.9701 1.0098 0.9254
[…]
Chinese Evaluation Results
Alibaba Cloud API 0.8988 0.8473 0.8942 0.9275 1.0223 1.0493 0.9124 0.8887 0.9301
XTTS 0.6563 0.4951 0.7149 0.5348 0.4681 0.4945 0.6191 0.6595 0.5803
CosyVoice2 0.7600 0.7739 0.7190 0.7943 0.7388 0.8843 0.6587 0.6597 0.7486
E2 TTS 0.6866 0.6513 0.7171 0.7082 0.6776 0.7027 0.6787 0.7161 0.6923
F5-TTS 0.7285 0.6901 0.7658 0.8052 0.8392 0.9291 0.6969 0.6890 0.7680
FireRedTTS 0.7536 0.6932 0.6968 0.7221 0.8371 0.8235 0.7093 0.6516 0.7359
Spark-TTS 0.7591 0.7179 0.7869 0.7887 0.8930 0.7902 0.7268 0.6890 0.7690
GPT-SoVITS 0.6915 0.6345 0.7269 0.6119 0.6583 0.6515 0.6296 0.6277 0.6540
MaskGCT 0.7391 0.5913 0.6558 0.7349 0.8011 0.8109 0.7722 0.6433 0.7186
[…]
Table 2: Comprehensive evaluation results of $1$ SRPAs on the SpeechRole-Eval benchmark. The table reports model performance across eight evaluation metrics: IA (Instruction Adherence), SF (Speech Fluency), $C$ (Conversational Coherence), SN (Speech Naturalness), PC (Prosody Consistency), EA (Emotion Appropriateness), PeC (Personality Consistency), and KC (Knowledge Consistency), along with the overall average score. Results are presented for both English and Chinese test sets. End-to-end models are indicated with a gray background. Proprietary models include Alibaba Cloud API and GPT-4o Audio, while all others are open-source.
5 Experiments
This section details our experimental setup and main results. We evaluate SRPAs under two paradigms: cascaded systems, which separate role-playing into ASR, LLM-based reasoning, and TTS stages, and end-to-end models that generate speech directly from prompts. All models are tested on the SpeechRole-Eval benchmark (Section 4) covering eight core metrics, with speaker similarity assessed for voice-controllable models. We also include supplementary experiments, such as fine-tuning a SRPA on our dataset to assess its speech role-playing performance.
5.1 SRPA Architectures
The primary objective of SRPAs is to generate coherent, spoken responses that are consistent with the target character’s identity upon receiving a user’s spoken query. To explore this, we examine two representative system architectures: the cascaded and end-to-end approaches.
Cascaded SRPAs decompose the overall task into three modular components. First, an ASR model transcribes the user’s spoken input into text. Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role. Finally, a TTS module converts the generated text into speech that matches the target character’s voice. In our implementation, both proprietary and open-source cascaded systems are explored. We utilize a proprietary cascaded system based on the Alibaba Cloud API [alibabacloud], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2. For the open-source side, we employ whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23] for high-quality ASR, LLaMA3.1-8B [DBLP:journals/corr/abs-2407-21783] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], E2 TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], F5-TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], GPT-SoVITS [gpt-sovits], Spark-TTS [DBLP:journals/corr/abs-2503-01710], and MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025].
End-to-end SRPAs generate speech responses directly from spoken input, guided by the user’s spoken query and a textual prompt of the character’s identity. Both the proprietary gpt-4o-audio-preview-2024-12-17 [DBLP:journals/corr/abs-2410-21276] and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] are evaluated in our study.
5.2 Evaluation Results
Main Evaluation Results.
We report the main evaluation results of all SRPAs on the SpeechRole-Eval benchmark. This benchmark covers eight core metrics: Instruction Adherence (IA), Speech Fluency (SF), Conversational Coherence (CC), Speech Naturalness (SN), Prosody Consistency (PC), Emotion Appropriateness (EA), Personality Consistency (PeC), and Knowledge Consistency (KC), with the final score computed as the average across these dimensions.
As shown in Table 2, the cascaded agents generally outperform end-to-end systems across most evaluation dimensions. Notably, models such as E2 TTS, F5-TTS, and Spark-TTS achieve relatively high overall scores in both content relevance and expressiveness, benefiting greatly from their modular design and optimization. While end-to-end models, particularly GPT-4o Audio, demonstrate competitive performance in expressiveness and coherence, they still exhibit noticeable shortcomings in style consistency and character setting. The performance of Qwen2.5-Omni is significantly lower across all metrics, indicating that end-to-end modeling remains challenging in the absence of sufficient task-specific supervision.
Language-Level Observations.
We observe significant differences between the English and Chinese subsets. Overall, the model performs better on English, particularly with respect to metrics such as content coherence and knowledge consistency. This disparity likely reflects the model’s language imbalance in pre-training and TTS capabilities. For example, XTTS performs relatively well in English but shows a marked decline in Chinese, especially in the expression of emotion and character personality. In contrast, Alibaba Cloud API and CosyVoice2 demonstrate more stable performance across both languages. These results indicate that current SRPAs exhibit uneven performance across different languages, with significantly better results in English than in Chinese. Therefore, further efforts are necessary to develop high-quality multilingual SRPAs.
Figure 4: Speaker similarity scores of SRPAs with voice cloning capability.
Speaker Similarity Comparison.
We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity. Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems. Following the same evaluation protocol as Seed-TTS [DBLP:journals/corr/abs-2406-02430], we use the WavLM-large [DBLP:journals/jstsp/ChenWCWLCLKYXWZ22] model to extract speaker embeddings and compute cosine similarity scores.
As shown in Figure 4, in the English setting, E2 TTS and F5-TTS achieve the highest speaker similarity, while GPT-SoVITS performs the worst. The overall ranking of models remains largely consistent across the two languages, indicating stable voice cloning performance. Notably, MaskGCT slightly outperforms others in Chinese, reaching the highest SIM score.
Models Evaluation Metrics Overall
IA SF $C$ SN PC EA PeC KC
English Evaluation Results
Qwen2.5-Omni 0.3977 0.4968 0.5491 0.4505 0.2958 0.3814 0.4089 0.4917 0.4340
SpeechRole-Agent 0.8072 0.6790 0.9480 0.6123 0.4202 0.5383 0.7568 0.9169 0.7098
[…]
Chinese Evaluation Results
Qwen2.5-Omni 0.4924 0.6886 0.5832 0.6204 0.4693 0.5125 0.3932 0.4772 0.5296
SpeechRole-Agent 0.9188 0.8233 0.9037 0.7749 0.7277 0.8607 0.8402 0.9253 0.8468
[…]
Table 3: Performance comparison between Qwen2.5-Omni and the fine-tuned SpeechRole-Agent on the SpeechRole-Eval benchmark. Results are reported for both English and Chinese, covering in-domain and out-of-domain scenarios. Out-of-domain results are indicated with gray background.
5.3 Fine-Tuning SRPA
Training Setup.
We fine-tune Qwen2.5-Omni-7B on the SpeechRole training set, containing 78 roles with speech-based dialogues. The model trains using speech inputs paired with role-consistent responses, enabling it to better understand and generate role-based spoken interactions. During evaluation, we use SpeechRole-Eval, dividing it into two subsets: in-domain, which includes roles seen during training, and out-of-domain, consisting of unseen roles. Results are reported separately for English and Chinese.
Performance Analysis.
After fine-tuning on the complete SpeechRole-Data training set, SpeechRole-Agent achieves notable gains over the original Qwen2.5-Omni. In the in-domain test set, the overall score in English increases from 0.4340 to 0.7098, while in Chinese it rises from 0.5296 to 0.8468. All evaluation metrics improve, with particularly notable advancements in instruction adherence, emotional appropriateness, and personality consistency. In the out-of-domain test set, the model also demonstrates strong generalization ability, with the English score increasing from 0.4786 to 0.6795 and the Chinese score from 0.5117 to 0.8028. These results fully demonstrate the model’s robust role-playing capabilities beyond the scope of the training set. Detailed results are presented in Table 3.
6 Conclusion
We propose SpeechRole, a unified framework for developing and evaluating SRPAs. To support this objective, we construct SpeechRole-Data, a large-scale, high-quality dataset comprising 112k dialogues from 98 characters, each with rich role settings and diverse vocal characteristics. Our benchmark establishes rigorous evaluation protocols for role-consistent speech generation and introduces SpeechRole-Eval, an automated multimodal evaluation suite that assesses various aspects including interaction competence, speech expressiveness, and role consistency.
Experimental results indicate current SRPAs still face challenges in achieving fully coherent and expressive role simulation, underscoring the need for more advanced character-based speech behavior modeling. We expect this work will provide a solid foundation for future research at the intersection of speech generation, role-playing, and embodied agents. All data, evaluation protocols, and baseline methods will be made publicly available to facilitate further studies on speech-interactive agents.
Figure 5: Left: Distribution of eight sub-tasks across three major task categories in SpeechRole. Right: Distribution of the 98 roles in SpeechRole-Data by language, gender, and data source.
Figure 6: Main evaluation results of $1$ SRPAs on SpeechRole-Eval. Left: English results. Right: Chinese results.Each metric is displayed with an interval range of 0.26, with all evaluation results normalized such that the maximum value across metrics is scaled to 1.30.
Appendix A SpeechRole-Data Construction
A.1 Text Data Construction
Role Selection.
A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024]. Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM. The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5). To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics. Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
Role Metadata Extraction.
For each character, we extract structured metadata, including: (1) Role Profile, such as temperament, behavioral tendencies, values, and personal preferences; (2) Background, including character settings, social identity, relationships, and living environment; and (3) Character Lines, including narration, monologues, and dialogue excerpts involving the target character in games or audiovisual works. These elements collectively define distinctive character profiles and support vivid, context-rich speech role-playing. Additionally, the dialogue excerpts serve as the basis for guiding dialogue generation and evaluating character consistency.
Dialogue Generation.
We use gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate dialogues between users and roles based on role metadata. For some roles with lengthy scripts, the content is divided into segments and processed sequentially. GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions. For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
Dialogue Filtering.
The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity. First, only dialogues that conform to a clear turn-taking pattern are retained. In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns. To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts. For English data, the all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] model is used, and samples with a similarity score below 0.4 are excluded. The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [text2vec] model, with a similarity score threshold of 0.45.
To further eliminate semantic redundancy, which is commonly observed in GPT-4.1-generated dialogues, near-duplicate samples are removed based on the similarity of their dialogue embeddings. Thresholds are set to 0.85 for Chinese and 0.9 for English. This de-duplication step is conducted separately for single-turn and multi-turn dialogues for each role. After filtering, the dataset contains 53,902 single-turn dialogues and 57,944 multi-turn dialogues.
Metrics IA SF $C$ SN PC EA PeC KC Overall
MSE 0.1020 0.0413 0.1426 0.0581 0.2539 0.1262 0.1690 0.1147 0.1260
Pearson 0.7304 0.6366 0.6076 0.8093 0.6316 0.8881 0.6677 0.7581 0.7162
Table 4: Evaluation results of correlation between automatic and human scoring.
A.2 Role Voice Collection
Role Voice Extraction.
To achieve authentic voice role-playing, real-life performance audio for each character is collected as reference material. For characters from Genshin Impact, in-game character audio is collected with permission from miHoYo. For other characters, audio is obtained from actual film and television works. Films and television series featuring these characters are gathered, and a total of 105 video files with an approximate duration of 181 hours are downloaded from the internet, covering all target characters.
During the preprocessing stage, all extracted speech segments undergo anonymization and fragmentation to ensure that they contain only brief, decontextualized utterances, with no retention of the original video or narrative context. As a result, no identifiable or copyrighted content remains, thus mitigating potential intellectual property risks and ensuring adherence to the principles of fair use for research purposes.
We first employ a batch processing pipeline based on ffmpeg [ffmpeg] to extract the raw audio streams from the collected video files. To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format. However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
Role Voice Processing.
To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [DBLP:conf/slt/HeSWLGHLYLSWCZW24] framework. This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation. The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22] metric. In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3. The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
Speaker Identification.
The extracted audio data contains only the speaker’s index and does not include identity information. To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437] models to determine the target character’s speaker number based on all transcribed texts. We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
Role Voice Ranking.
To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity. Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [DBLP:conf/interspeech/WangZCC023] model and calculate the cosine similarity between each sample and all other samples. A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role. Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.
Appendix B Human Correlation Verification Details
To evaluate the consistency between automatic scoring and human scoring, we randomly selected 20 samples from the dataset, including 5 single-turn dialogue samples and 15 multi-turn dialogue samples. For each sample, we randomly selected two model-generated speech responses for evaluation.
Gemini scored each model response and its corresponding reference response on a scale of 1 to 10 based on 8 predefined scoring metrics. The final score for a model was calculated as the ratio of the model response score to the reference response score.
For human evaluation, four trained evaluators were employed to directly compare the responses of the two models based on the same 8 scoring metrics. The evaluation criteria were as follows: the superior model received 1 point, the inferior model 0 points, and if the performance was deemed equivalent, each received 0.5 points. The final human score was determined by averaging the scores from all evaluators.
To quantify the agreement between automated and human evaluation methods, we computed two key metrics: mean squared error (MSE) and Pearson correlation coefficient. The MSE was calculated by first fitting a linear regression between the score differences of paired model responses from human evaluators and gemini. The predicted gemini score differences were then derived from this regression model, and the MSE was computed between these predicted values and the actual gemini score differences. For the Pearson correlation analysis, we computed the linear correlation coefficient between the human evaluators’ score differences and gemini’s score differences across all paired comparisons.
Table 4 reports the detailed results, the overall MSE of 0.1260 indicates relatively minor deviations between the two scoring methods. The high Pearson correlation coefficient further confirms the reliability of the automated scoring system. The evaluation results collectively demonstrate strong agreement between automatic and human scoring across all metrics, indicating that the automated scoring system can effectively approximate human judgment.
Appendix C Additional data statistics
Role Type Distribution
Figure 5 presents the distribution of characters by language (Chinese and English), source type (game, film, television series, and animation), and gender. English characters, particularly those from films, constitute the majority in the dataset. In contrast, the Chinese subset exhibits a more balanced distribution across various source types. This diversity provides extensive coverage of voices and contexts for role-playing.
Task Type Distribution
User prompts can generally be categorized into three major task categories, which can be further divided into eight subtask types, as illustrated in the figure 5. The description and proportion of each subtask type are as follows:
  1. 1.
    Internal Reasoning
    • Opinion and Emotion Inquiry: Questions about opinions, emotions, inner feelings, or reactions. (14.73%)
    • Reasoning and Motivation: Questions about why someone did something, their motivations, or decision-making processes. (14.73%)
    • Reflection and Change: Questions about how practices/opinions have changed. (9.55%)
  2. 2.
    Experiential Narration
    • Action Description and Retrospection: Questions about what happened, how something was done, or detailed retrospectives. (14.72%)
    • Skills and Abilities: Questions about how to accomplish something, difficulties with skills, or ability changes. (6.38%)
    • Historical and Background Description: Questions about history or background of places/organizations/people. (11.55%)
  3. 3.
    Social Communication
    • Judgment of Others and Events: Questions about relationships, event evaluations, or interpersonal interactions. (14.73%)
    • Advice and Life Experience: Questions seeking advice, coping methods, or life experiences. (13.61%)
Evaluation Results Demonstration
Figure 6 visually presents the evaluation results of all SRPAs on the SpeechRole-Eval benchmark.
Appendix D Prompt Templates
Prompts for Batch Data Generation
Figure 7 and Figure 8 respectively present the prompts used for generating single-turn dialogue data for English and Chinese characters. Figure 9 and Figure 10 respectively present the prompts used for generating multi-turn dialogue data for English and Chinese characters.
Prompts for Automated Judgement
Figure 11 displays the judgment prompt designed for automatically assessing audio responses generated by the two models.
Appendix E SpeechRole-Agent training settings
Figure 12 presents the distributed training script for fine-tuning the Qwen2.5-Omni-7B model on 8×H20 GPUs. The configuration employs bfloat16 mixed-precision training with a learning rate of 1e-4 and gradient accumulation. The implementation includes periodic evaluation (every 500 steps) and checkpoint management.
Role Name Category Language Gender Source Split
hutao Game Chinese Female Genshin Impact train/test
raidenShogun Game Chinese Female Genshin Impact train/test
wanderer Game Chinese Male Genshin Impact train/test
ayaka Game Chinese Female Genshin Impact dev/test
zhongli Game Chinese Male Genshin Impact train/test
liyunlong TV Chinese Male Drawing Sword dev/test
wangduoyu Movie Chinese Male Hello Mr. Billionaire train/test
weixiaobao TV Chinese Male The Deer and the Cauldron train/test
jiumozhi TV Chinese Male Demi-Gods and Semi-Devils train/test
wangyuyan TV Chinese Female Demi-Gods and Semi-Devils train/test
Luna Movie English Female Harry Potter dev/test
Penny TV English Female The Big Bang Theory dev/test
zhangwuji TV Chinese Male The Heaven Sword and Dragon Saber train/test
zhaomin TV Chinese Female The Heaven Sword and Dragon Saber train/test
huangrong TV Chinese Female The Legend of the Condor Heroes train/test
guojing TV Chinese Male The Legend of the Condor Heroes dev/test
wukong TV Chinese Male Journey to the West train/test
HAL 9000 Movie English Male 2001: A Space Odyssey train/test
Colonel Nathan R. Jessep Movie English Male A Few Good Men train/test
Antonio Salieri Movie English Male Amadeus train/test
Stifler Movie English Male American Pie train/test
Paul Vitti Movie English Male Analyze That train/test
Alvy Singer Movie English Male Annie Hall train/test
Violet Weston Movie English Female August: Osage County train/test
Willie Soke Movie English Male Bad Santa train/test
Gaston Animation English Male Beauty and the Beast train/test
The Dude Movie English Male The Big Lebowski train/test
Paul Conroy Movie English Male Buried train/test
Truman Capote Movie English Male Capote train/test
Mater Animation English Male Cars 2 train/test
Andrew Detmer Movie English Male Chronicle train/test
Coriolanus Movie English Male Coriolanus train/test
John Keating Movie English Male Dead Poets Society dev/test
Wade Wilson Movie English Male Deadpool dev/test
Jim Morrison Movie English Male The Doors train/test
Queen Elizabeth I Movie English Female Elizabeth: The Golden Age dev/test
Jeff Spicoli Movie English Male Fast Times at Ridgemont High train/test
Fred Flintstone Animation English Male The Flintstones train/test
Freddy Krueger Movie English Male Freddy Vs.Jason train/test
Tyrion Lannister TV English Male Game of Thrones train/test
James Brown Movie English Male Get on Up train/test
Walt Kowalski Movie English Male Gran Torino train/test
John Coffey Movie English Male The Green Mile train/test
Theodore Twombly Movie English Male Her dev/test
Gregory House TV English Male House M.D. dev/test
Sonny Movie English Male I, Robot train/test
Colonel Hans Landa Movie English Male Inglourious Basterds train/test
Judge Dredd Movie English Male Judge Dredd dev/test
Juno MacGuff Movie English Female Juno train/test
Professor G.H. Dorr Movie English Male The Ladykillers train/test
Table 5: Summary of character attributes and dataset partitioning (part 1).
Role Name Category Language Gender Source Split
Fletcher Reede Movie English Male Liar Liar train/test
Abraham Lincoln Movie English Male Lincoln train/test
Frank T.J. Mackey Movie English Male Magnolia train/test
Leonard Shelby Movie English Male Memento train/test
Harvey Milk Movie English Male Milk train/test
Randle McMurphy Movie English Male One Flew Over the Cuckoo’s Nest train/test
Jack Sparrow Movie English Male Pirates of the Caribbean: Dead Man’s Chest dev/test
John Dillinger Movie English Male Public Enemies train/test
Lestat de Lioncourt Movie English Male The Queen of Damned train/test
Tyler Hawkins Movie English Male Remember Me dev/test
James Carter Movie English Male Rush Hour 2 train/test
Jigsaw Movie English Male Saw train/test
John Doe Movie English Male Se7en train/test
Sherlock Holmes Movie English Male Sherlock Holmes dev/test
Shrek Animation English Male Shrek train/test
Pat Solitano Movie English Male Silver Linings Playbook train/test
Karl Childers Movie English Male Sling Blade train/test
Bruno Antony Movie English Male Strangers on a Train train/test
Seth Movie English Male Superbad train/test
Caden Cotard Movie English Male Synecdoche, New York train/test
Travis Bickle Movie English Male Taxi Driver train/test
Stanley Ipkiss Movie English Male The Mask dev/test
Lyn Cassady Movie English Male The Men Who Stare at Goats train/test
Michael Scott TV English Male The Office dev/test
Robert Angier Movie English Male The Prestige dev/test
Dr. Frank-N-Furter Movie English Male The Rocky Horror Picture Show train/test
Jack Torrance Movie English Male The Shining train/test
Tom Ripley Movie English Male The Talented Mr. Ripley train/test
D_Artagnan Movie English Male The Three Musketeers train/test
Thor Movie English Male Thor: Ragnarok train/test
James Bond Movie English Male Tomorrow Never Dies dev/test
Mark Renton Movie English Male Trainspotting train/test
David Aames Movie English Male Vanilla Sky train/test
Rorschach Movie English Male Watchmen train/test
Jordan Belfort Movie English Male The Wolf of Wall Street train/test
Logan Movie English Male X-Men Origins: Wolverine dev/test
Judy Hoops Animation English Female Zootopia train/test
Doctor Who TV English Male Doctor Who train/test
Raylan Givens TV English Male Justified train/test
Mary Sibley TV English Female Salem train/test
Lucifer Morningstar TV English Male Lucifer train/test
Twilight Sparkle Animation English Female My Little Pony: Friendship is Magic dev/test
Oliver Queen TV English Male Arrow train/test
Klaus Mikaelson TV English Male The Originals train/test
Queen Catherine TV English Female Reign train/test
Dr. Hannibal Lecter TV English Male Hannibal train/test
Coach Eric Taylor Movie English Male Friday Night Lights train/test
yaemiko Game Chinese Female Genshin Impact train/test
Table 6: Summary of character attributes and dataset partitioning (part 2).
Figure 7: Prompt structure for batch single-turn English dialogue generation.
Figure 8: Prompt structure for batch single-turn Chinese dialogue generation.
Figure 9: Prompt structure for batch multi-turn English dialogue generation.
Figure 10: Prompt structure for batch multi-turn Chinese dialogue generation.
Figure 11: Prompt structure for Automated Judgement.
Figure 12: Training script configuration for fine-tuning the Qwen2.5-Omni-7B model.

### v6

@@
 Abstract
-Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance.
-Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios.
-In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs).
-To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations.
-Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing.
-Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity.
-Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence.
-We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
-Abstract
-Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance.
-Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios.
-In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs).
-To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations.
-Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing.
-Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity.
-Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence.
-We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
-[…]
-[…]
+Speech is essential for realistic role-playing, yet existing work on role-playing agents largely centers on text, leaving Speech Role-Playing Agents (SRPAs) underexplored and lacking systematic evaluation.
+We introduce SpeechRole, a unified framework for developing and assessing SRPAs.
+SpeechRole-Data contains 98 roles and 111k speech-to-speech conversations with rich timbre and prosodic variation, providing large-scale resources for training SRPAs.
+SpeechRole-Eval offers a multidimensional benchmark that directly evaluates generated speech, preserving paralinguistic cues and measuring interaction ability, speech expressiveness, and role-playing fidelity.
+Experiments show that end-to-end SRPAs such as GPT-4o Audio achieve strong fluency and naturalness, but remain limited in prosody consistency and emotion appropriateness.
+In contrast, current open-source end-to-end models exhibit substantial performance gaps across multiple evaluation dimensions.
+Cascaded and end-to-end systems achieve comparable results in interaction ability and role-playing fidelity, suggesting that these aspects are still largely influenced by the underlying text-based language models.
+We release all data, code, and evaluation tools at https://github.com/yuhui1038/SpeechRole.
+*Equal Contribution.
+Corresponding Author.
 Figure 1: Overview of the SpeechRole framework.
@@
 However, current research predominantly focuses on the text modality [DBLP:journals/corr/abs-2308-09597, DBLP:conf/emnlp/ShaoLDQ23, DBLP:conf/acl/WangPQLZWGGN00024, DBLP:conf/emnlp/SadeqXKLGM24], overlooking the crucial role of speech in authentic interactions.
+Speech Role-Playing Agents (SRPAs) are speech-to-speech systems designed to generate spoken responses in character, reproducing distinctive vocal timbre, prosody, and persona traits while maintaining coherence across interactions.
 Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [DBLP:conf/iui/ChangSZJ0K23].
-Although preliminary studies have explored role-playing incorporating visual or speech modalities [DBLP:conf/iclr/DaiH0J0025, DBLP:journals/corr/abs-2505-20277], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs).
-SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions.
-In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
-To address this gap, we propose SpeechRole, a unified framework for the development and evaluation of SRPAs (see Figure 1).
-We first introduce SpeechRole-Data, a large-scale dataset comprising 112k speech-based dialogues covering 98 diverse roles.
-Each role features distinctive timbre and prosodic characteristics, enabling expressive and role-consistent speech interactions.
-Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses agents across three dimensions: interaction competence, speech expressiveness, and role consistency.
-This benchmark supports both cascaded and end-to-end architectures, facilitating comprehensive and comparable analyses across different paradigms.
-Our experiments reveal complementary strengths and weaknesses between SRPA paradigms.
-Cascaded systems offer strong modular control and stable role identity across interactions, but may suffer from cross-module mismatches or limited expressiveness.
-In contrast, end-to-end models produce more coherent and natural speech, yet often struggle with speaker consistency and preserving character traits over extended dialogues.
-These trade-offs highlight the unique challenges in building effective SRPAs that balance interaction ability, speech expressiveness, and role-playing fidelity.
+Although preliminary studies have explored speech-based role-playing [DBLP:journals/corr/abs-2505-20277], large-scale datasets and systematic evaluation benchmarks for SRPAs remain scarce.
+To address this gap, we introduce SpeechRole, a unified framework for developing and evaluating SRPAs.
+SpeechRole-Data is a large-scale speech-to-speech dataset containing 111k dialogues across 98 roles, offering diverse vocal traits and prosodic patterns that support the construction of SRPAs.
+Building on this foundation, SpeechRole-Eval provides a multidimensional benchmark that directly evaluates generated speech without relying on speech-to-text conversion, thus preserving critical paralinguistic information.
+The benchmark assesses SRPAs along three complementary dimensions: interaction ability, speech expressiveness, and role-playing fidelity.
+Together, SpeechRole-Data and SpeechRole-Eval establish a foundation for systematic research on speech-based role-playing.
+Our empirical findings highlight clear differences between cascaded and end-to-end SRPAs.
+The most advanced end-to-end system, GPT-4o Audio, demonstrates notable advantages in speech fluency and naturalness, while still exhibiting limitations in prosody consistency and emotion appropriateness.
+In contrast, current open-source end-to-end models show substantial performance gaps across multiple evaluation dimensions.
+Meanwhile, for interaction ability and role-playing fidelity, cascaded and end-to-end systems achieve comparable results, suggesting that these capabilities remain strongly influenced by the underlying text-based large language models rather than the specific speech modeling pipeline.
 Overall, our main contributions are as follows:
   1. 1.
-    We construct SpeechRole-Data, a large-scale multilingual corpus covering 98 roles and 112k speech-based dialogues with rich prosodic and timbral variation.
+    We construct SpeechRole-Data, a large-scale speech-to-speech corpus with 111k dialogues across 98 roles, offering diverse timbral and prosodic patterns for developing SRPAs.
   2. 2.
-    We develop SpeechRole-Eval, a comprehensive benchmark designed to evaluate SRPA capabilities across key dimensions including interaction ability, speech expressiveness, and role-playing fidelity.
+    We introduce SpeechRole-Eval, a multidimensional benchmark that directly evaluates generated speech and assesses interaction ability, speech expressiveness, and role-playing fidelity without relying on speech-to-text conversion.
   3. 3.
-    We systematically compare cascaded and end-to-end paradigms, revealing their strengths, limitations, and offering insights for future SRPA development.
+    We provide a systematic comparison of cascaded and end-to-end SRPAs, characterizing their strengths and limitations and analyzing how system design impacts speech quality and role-playing performance.
+Figure 2: Examples of the three major task categories in SpeechRole.
 2 Related Work
-Role-Playing Agents Across Modalities.
-Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality.
-ChatHaruhi [DBLP:journals/corr/abs-2308-09597] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], Character-LLM [DBLP:conf/emnlp/ShaoLDQ23], and RoleFact [DBLP:conf/emnlp/SadeqXKLGM24] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations.
-Beyond text, MMRole [DBLP:conf/iclr/DaiH0J0025] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the value of speech for real-time, voice-aware character interaction.
-To assess role-playing capabilities, benchmarks such as RoleEval [DBLP:journals/corr/abs-2312-16132], RoleInteract [DBLP:journals/corr/abs-2403-13679], SocialBench [DBLP:conf/acl/ChenCYXXSQLZH24], and CharacterEval [DBLP:conf/acl/TuFTSSGY24] evaluated behavioral consistency and persona fidelity, while INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24] and ECHO [DBLP:journals/corr/abs-2404-13957] introduced human-aligned, psychometric evaluation paradigms.
-These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity.
-However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
-Interactive Speech Agents.
-In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end.
+2.0.0.1 Role-Playing Agents Across Modalities.
+Recent research has increasingly explored the role-playing capabilities of LLMs, with most efforts focusing on the text modality.
+Notable examples include ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], which propose methods to improve persona simulation and reduce hallucinations.
+OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the importance of incorporating speech for real-time, voice-aware character interactions.
+Existing benchmarks, such as RoleEval [DBLP:journals/corr/abs-2312-16132] and INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24], focus on behavioral consistency and persona fidelity.
+However, systematic evaluation frameworks for speech-based role-playing remain scarce, leaving a gap in assessing vocal expression, emotional depth, and user engagement.
+2.0.0.2 Interactive Speech Agents.
+Speech-interactive LLMs typically follow two paradigms: cascaded and end-to-end.
 Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS).
-This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24].
-Modern TTS models like F5TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], E2-TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], Spark-TTS [DBLP:journals/corr/abs-2503-01710], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025], and XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24] have significantly improved output quality, expressiveness, and controllability.
-However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
+This modular design has been widely adopted in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24].
+Recent TTS models have improved naturalness, expressiveness, and controllability, although cascaded pipelines can suffer from error accumulation and limited cross-modal context integration.
 End-to-end systems aim to unify speech perception and generation within a single model.
-Early efforts like SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [DBLP:journals/corr/abs-2306-12925] utilized vocoder-based representations for continuous speech modeling.
-Recent advances, including Moshi [DBLP:journals/corr/abs-2410-00037], LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025], and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
-Figure 2: Examples of the three major task categories in SpeechRole.
+Early work such as SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies.
+More recent architectures, including LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025] and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], advance end-to-end speech modeling through dual-modality training, streaming capabilities, and low-latency generation, enabling more natural and efficient speech-driven interactions.
+Despite these advances, systematic evaluation of speech role-playing capabilities in end-to-end settings remains limited.
 3 SpeechRole-Data
-We introduce SpeechRole-Data, a large-scale dataset for speech-based role-play, featuring 98 roles and over 112k dialogues with reference audio and rich character profiles.
-To ensure personality and voice consistency, we follow a structured pipeline spanning role selection, dialogue generation, and audio processing.
-Details are in the appendix.
+SpeechRole-Data is constructed to support SRPAs, which require speech-based dialogues that capture not only a character’s linguistic behavior but also their distinctive vocal attributes.
+Building such data necessitates two essential components: persona-consistent dialogue content that reflects how the character would interact, and representative voice samples that preserve their timbre, prosody, and speaking style.
+To obtain both elements, we curate 98 characters drawn from television dramas, films, animations, and games.
+For each character, we collect narrative profiles and scripts.
+Based on these materials, an LLM is prompted to generate user–character dialogue texts that remain faithful to the original personality and storyline.
+In parallel, we extract clean speech segments from the corresponding audiovisual sources to capture each character’s timbre and prosody.
+Using the generated texts and collected voice samples, we synthesize full speech-to-speech dialogues with a TTS system.
+This pipeline produces a corpus of 111k spoken interactions tailored for SRPA development and evaluation.
+Splits Characters Samples (single-turn/multi-turn) Speech Hours (user/character)
+Train 78 89,461 (43,082/46,379) 875.25 (204.45/670.80)
+Out-of-domain 20 21,993 (10,627/11,366) 203.60 (49.84/153.76)
+Test 98 392 (196/196) 4.20 (0.94/3.26)
+Table 1: Statistics of SpeechRole-Data across train, dev, and test splits.
 3.1 Text Data Construction
-Role Selection.
+3.1.0.1 Role Selection.
 We curate 98 diverse roles, including 18 from ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and 80 from RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024].
 The selection balances language, gender, and personality, focusing on characters with distinctive vocal traits to enrich speech modeling.
-Role Metadata Extraction.
+3.1.0.2 Role Metadata Extraction.
 For each character, we extract structured metadata to build a comprehensive profile, including (1) Role Profile (temperament, preferences), (2) Background (social identity, relationships), and (3) Character Lines (dialogues, monologues).
 This metadata forms the basis for generating personality-consistent dialogues.
-Dialogue Generation.
+3.1.0.3 Dialogue Generation.
 Using the extracted metadata as a guide, we employ gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate conversations.
 For each of the 98 roles, we produce approximately 800 single-turn and 800 multi-turn dialogues, creating a substantial initial pool of text data.
-Dialogue Filtering and Deduplication.
-The generated data undergoes a multi-stage filtering process.
-First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese.
-To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
-Splits Characters Samples (single-turn/multi-turn) Speech Hours (user/character)
-Train 78 89,461 (43,082/46,379) 875.25 (204.45/670.80)
-Dev 20 21,993 (10,627/11,366) 203.60 (49.84/153.76)
-Test 98 392 (196/196) 4.20 (0.94/3.26)
-Table 1: Statistics of SpeechRole-Data across train, dev, and test splits.
-3.2 Role Voice Collection
-Voice Collection and Preprocessing.
-To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works.
-All audio is anonymized and fragmented into brief utterances to comply with fair use.
+3.1.0.4 Dialogue Deduplication.
+To improve data diversity and avoid repeated conversational patterns, we perform dialogue-level deduplication using semantic similarity.
+For each dialogue, we compute its embedding-based similarity with all other dialogues using all-MiniLM-L6-v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese.
+Dialogues with similarity above 0.9 for English or 0.85 for Chinese are considered near-duplicates, and only one instance is retained.
+3.2 Role Voice Collection and Synthesis
+3.2.0.1 Voice Collection and Preprocessing.
+To acquire authentic reference voices, we collect audio from sources such as the game Genshin Impact, whose assets permit non-commercial academic research use.
+For other audiovisual works, we only extract short speech fragments for analysis and dataset construction and do not distribute or reproduce any original copyrighted material.
+All collected audio is anonymized and segmented into brief utterances.
 Raw streams are extracted using ffmpeg [ffmpeg] and converted to mono 16 kHz WAV format.
-Voice Cleaning and Annotation.
-We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, including source separation, speaker diarization, and voice activity detection.
-This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23] and scored with DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22].
-Figure 3: Speech duration distribution of users and roles.
-Speaker Identification.
-The diarization process outputs speaker indices without explicit identity information.
-To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to analyze transcribed text.
-The models’ integrated results are then manually verified.
-Reference Audio Ranking and Selection.
-To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity.
-We use the CAM++ model [DBLP:conf/interspeech/WangZCC023] to extract a speaker embedding for each clip.
-The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
+3.2.0.2 Audio Cleaning and Segmentation.
+We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, performing source separation, speaker diarization, and voice activity detection.
+This pipeline produces clean 3–10 second single-speaker clips.
+The quality of each clip is assessed using DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22], and clips with an overall quality score below 3 are filtered out.
+3.2.0.3 Role-Level Speaker Identification.
+Speaker diarization groups the original audiovisual audio into segments corresponding to different speakers, assigning each speaker a numeric ID but not revealing which character the ID represents.
+To map each segment to the correct character, we first transcribe the clips using whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23].
+We then use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to infer character identity based on the transcribed text and contextual cues.
+Finally, all assignments are manually verified for accuracy.
+3.2.0.4 Reference Voice Selection.
+We select a representative reference voice for each role by ranking all its cleaned clips according to vocal consistency.
+For each clip, we extract a speaker embedding using the CAM++ model [DBLP:conf/interspeech/WangZCC023].
+We then compute its average cosine similarity to all other clips belonging to the same role and choose the clip with the highest similarity as the reference voice, as it best reflects the role’s characteristic timbre and speaking style.
+3.2.0.5 Role and User Speech Synthesis.
+Using the collected reference voices, we synthesize role utterances with CosyVoice2 [DBLP:journals/corr/abs-2412-10117], which conditions on each role’s selected clip to preserve timbre and prosody.
+User utterances are generated with the VolcEngine TTS system [volcengine] using a single fixed voice, as the user does not require persona-specific vocal variation.
+This yields complete speech-to-speech dialogues aligned with both the textual content and role-specific vocal characteristics.
+Figure 3: Speech duration distribution of SpeechRole-Data.
 3.3 Dataset Statistics
-Dataset Overview.
-SpeechRole-Data comprises 98 distinct roles and 112k dialogue samples, including both single-turn and multi-turn conversations (see Table 1).
-Each role is accompanied by a detailed role profile and a reference audio clip that demonstrates its unique speaking style, such as timbre, rhythm, and prosody.
-Each dialogue sample contains both text and speech, along with contextual background information relevant to the role, thereby supporting realistic, evidence-based role-playing.
-During the evaluation phase, the test set consists of two single-turn and two multi-turn dialogues sampled for each role.
-The remaining data are divided into a training set (78 roles) and a validation set (20 roles).
+3.3.0.1 Dataset Overview.
+SpeechRole-Data comprises 98 distinct roles and 111k speech-to-speech dialogue samples, covering both single-turn and multi-turn conversations.
+Each role is paired with a detailed role profile and a reference audio clip that illustrates its characteristic speaking style, including timbre, rhythm, and prosody.
+Each dialogue sample contains speech data and contextual information relevant to the role, enabling realistic and evidence-grounded role-playing.
+The dataset is divided into three subsets.
+The training set contains dialogues from 78 roles.
+The out-of-domain set includes 20 roles that do not appear in the training data, enabling evaluation of generalization to unseen characters.
+The test set covers all 98 roles, with four carefully curated dialogue samples per role, resulting in a total of 392 test instances.
+These instances are independently selected from both the training and out-of-domain pools (see Table 1).
+The scale of this test set is comparable to prior role-playing evaluation benchmarks, such as the 400-sample evaluation in DBLP:journals/corr/abs-2505-20277 and the 294-sample benchmark introduced by DBLP:conf/iclr/DaiH0J0025.
 Detailed examples are provided in Figure 2.
-Speech Duration.
-We analyze the speech duration of users and roles.
-As shown in Figure 3, the duration of role utterances is generally longer, typically ranging from 10 to 20 seconds.
-This is mainly due to the narrative or expressive nature of their content.
-In contrast, user utterances are concentrated within 3 to 6 seconds, reflecting their concise and goal-oriented characteristics.
+3.3.0.2 Speech Duration.
+We analyze the duration of individual speech segments for both users and roles.
+As shown in Figure 3, single utterances from roles are generally longer, typically ranging from 10 to 20 seconds, reflecting their narrative or expressive nature.
+In contrast, user utterances are concentrated between 3 and 6 seconds, highlighting their concise and goal-oriented characteristics.
+Models IA $C$ SF SN PC EA PeC KC Overall
+English Evaluation Results
+Qwen3-8B 0.926 0.955 0.861 0.825 0.845 0.884 0.871 0.908 0.884
+Llama-3.1-8B 0.915 0.932 0.846 0.823 0.824 0.864 0.826 0.886 0.865
+Mistral-7B 0.843 0.894 0.859 0.820 0.781 0.806 0.752 0.819 0.822
+Alibaba Cloud API 0.895 0.954 0.823 0.786 0.768 0.799 0.784 0.867 0.835
+LLaMA-Omni 0.791 0.865 0.880 0.793 0.674 0.668 0.659 0.734 0.758
+Qwen2.5-Omni 0.537 0.647 0.682 0.640 0.464 0.490 0.403 0.595 0.557
+GPT-4o Audio 0.901 0.990 0.942 0.912 0.811 0.846 0.807 0.881 0.886
+Chinese Evaluation Results
+Qwen3-8B 0.823 0.871 0.858 0.829 0.874 0.875 0.844 0.858 0.854
+Llama-3.1-8B 0.818 0.776 0.820 0.822 0.785 0.798 0.714 0.748 0.785
+Mistral-7B 0.717 0.700 0.840 0.811 0.712 0.702 0.602 0.664 0.719
+Alibaba Cloud API 0.905 0.929 0.919 0.913 0.861 0.865 0.828 0.827 0.881
+Qwen2.5-Omni 0.644 0.767 0.890 0.799 0.608 0.620 0.511 0.652 0.686
+GPT-4o Audio 0.858 0.961 0.902 0.896 0.828 0.841 0.800 0.868 0.869
+Table 2: Comprehensive evaluation results of 7 SRPAs on the SpeechRole-Eval benchmark.
+The table reports model performance across eight evaluation metrics: IA (Instruction Adherence), $C$ (Conversational Coherence), SF (Speech Fluency), SN (Speech Naturalness), PC (Prosody Consistency), EA (Emotion Appropriateness), PeC (Personality Consistency), and KC (Knowledge Consistency), along with the overall average score.
+Results are presented for both English and Chinese test sets.
+End-to-end models are indicated with a gray background.
+Proprietary models include Alibaba Cloud API and GPT-4o Audio, while all others are open-source.
 4 SpeechRole-Eval
-To systematically assess SRPAs, we introduce SpeechRole-Eval, a comprehensive benchmark that evaluates model performance across three key dimensions: fundamental interaction ability, speech expressiveness, and role-playing fidelity.
-In contrast to prior work that primarily focuses on text, SpeechRole-Eval places special emphasis on the speech modality as a critical component of evaluation.
-The scoring process is fully automated, powered by a strong multimodal model, and has been validated against human judgments to ensure reliability.
+To systematically assess SRPAs, we introduce SpeechRole-Eval, a benchmark designed to evaluate model performance across three key dimensions: interaction ability, speech expressiveness, and role-playing fidelity.
+Existing evaluations of speech-based role-playing remain largely reliant on human judgments [DBLP:journals/corr/abs-2505-20277], lacking standardized, objective metrics.
+At the same time, recent research has demonstrated the growing potential of LLMs as automated judges in various evaluation tasks [li2025generation].
+Motivated by these developments, SpeechRole-Eval leverages a LLM with speech understanding capabilities to provide automated, objective, and reliable assessment of SRPAs, with the scoring process validated against human annotations.
 This section describes the task setup, evaluation metrics, and the overall evaluation pipeline.
-4.1 Task Types
-To assess the cognitive, affective, and social competencies of SRPAs, user prompts are categorized into three major task categories based on communicative intent and interactional focus: (1) Internal Reasoning, which concerns the character’s rich subjective world, including personal perspectives, emotional responses, and underlying motivations; (2) Experiential Narration, which emphasizes the character’s ability to vividly recall past events, describe concrete actions, and provide relevant background or historical information; and (3) Social Communication, which highlights interpersonal understanding achieved through nuanced evaluative judgments, the provision of advice, and relational reasoning in context.
-These categories reflect the core dimensions of role-playing dialogue and enable a comprehensive, structured assessment of role-playing abilities (see Figure 2).
 4.1 Evaluation Metrics
-In contrast to traditional text-based RPAs, SRPAs must not only accurately emulate specific characters but also consistently deliver coherent and expressive spoken responses fully aligned with both the dialogue context and the character’s unique persona.
-To comprehensively evaluate SRPAs, we propose a robust three-dimensional evaluation framework that assesses fundamental interaction ability, speech quality and expressiveness, and role-playing fidelity.
-Fundamental interaction abilities comprehensively evaluate the agent’s capacity to maintain coherent, natural, and instruction-aligned interactions fully consistent with the character, which are assessed by the following key metrics:
+SRPAs must handle not only the linguistic demands of dialogue generation but also the vocal and stylistic requirements of spoken character portrayal.
+Evaluating such systems therefore requires attention to aspects beyond those considered in traditional text-based settings.
+To capture the full range of SRPA capabilities, we adopt a three-dimensional evaluation framework.
+Interaction Ability: measures the agent’s capacity to produce coherent, contextually appropriate, and instruction-aligned responses.
+Assessed using Instruction Adherence (IA) and Conversational Coherence (CC).
+Speech Quality and Expressiveness: evaluates the naturalness, fluency, and prosodic expressiveness of speech.
+Assessed using Speech Fluency (SF), Speech Naturalness (SN), Prosodic Consistency (PC), and Emotion Appropriateness (EA).
+Role-Playing Fidelity: examines how well the agent maintains character-specific personality and knowledge. Assessed using Personality Consistency (PeC) and Knowledge Consistency (KC).
+4.2 LLM-Based Evaluation Procedure
+There is a lack of widely recognized objective metrics for evaluating role-conditioned speech outputs, and absolute scoring by LLMs is known to be unstable and inconsistent, largely due to the open-ended nature of role-playing, the absence of a single correct response, and the sensitivity of absolute scores to prompt phrasing and scoring scale [DBLP:conf/acl/WangLCCZLCKLLS24].
+To address these limitations, recent work advocates pairwise scoring as a more reliable alternative [DBLP:conf/iclr/DaiH0J0025], as it bypasses the need for a ground-truth answer, improves inter-judge consistency, and produces preference signals that better align with human judgments.
+Motivated by these findings, SpeechRole-Eval adopts a pairwise relative scoring protocol using an LLM with strong speech understanding capabilities, enabling stable, objective, and reproducible evaluation of SRPAs.
+To enhance evaluation consistency and interpretability, the LLM is first prompted to produce a concise rationale before issuing its rating, encouraging more deliberate judgment.
+Moreover, instead of assessing each speech sample independently, we adopt a pairwise comparative setting, where the LLM jointly receives a test sample and a high-quality reference response and assigns scores to both, enabling more stable and discriminative relative assessments.
+We employ gemini-2.5-pro [DBLP:journals/corr/abs-2312-11805] to assign scores to each response following the established evaluation protocol, with ratings ranging from 1 to 10.
+For each test speech sample, the final evaluation score is computed as the ratio of its rating to that of a high-quality reference response.
+Formally, for $N$ evaluation instances (where $N=392$, corresponding to the size of our test set), the metric is defined as
+$\text{Score}=\frac{1}{N}\sum_{i=1}^{N}\frac{s_{i}^{\text{test}}}{s_{i}^{\text{ref}}}.$
+This ratio standardizes scores across different prompts and reduces variance caused by prompt differences, thereby improving stability and promoting better alignment with human judgments.
+Models IA $C$ SF SN PC EA PeC KC Overall
+English Evaluation Results
+Qwen2.5-Omni 0.539 0.657 0.677 0.633 0.458 0.483 0.399 0.602 0.556
+SpeechRole-Agent 0.806 0.957 0.881 0.791 0.608 0.689 0.608 0.883 0.778
+Qwen2.5-Omni 0.531 0.608 0.700 0.666 0.489 0.518 0.417 0.567 0.562
+SpeechRole-Agent 0.720 0.940 0.864 0.751 0.566 0.625 0.557 0.795 0.727
+Chinese Evaluation Results
+Qwen2.5-Omni 0.654 0.766 0.891 0.822 0.615 0.629 0.525 0.640 0.693
+SpeechRole-Agent 0.908 0.935 0.896 0.849 0.763 0.800 0.739 0.859 0.843
+Qwen2.5-Omni 0.599 0.773 0.884 0.700 0.578 0.580 0.450 0.707 0.659
+SpeechRole-Agent 0.859 0.925 0.940 0.832 0.768 0.815 0.752 0.909 0.850
+Table 3: Performance comparison between Qwen2.5-Omni and the fine-tuned SpeechRole-Agent on the SpeechRole-Eval benchmark.
+Results are reported for both English and Chinese, covering in-domain and out-of-domain scenarios.
+Out-of-domain results are indicated with gray background.
+5 Experiments
+This section presents a comprehensive experimental evaluation of SRPAs on SpeechRole-Eval.
+We evaluate a diverse set of representative baselines spanning both cascaded and end-to-end system paradigms, which reflect the two dominant design choices for current SRPAs.
+Experiments are conducted in both Chinese and English settings to assess multilingual performance and generalization.
+In addition, we analyze the effect of role-specific training by comparing a base end-to-end model with its counterpart further trained on SpeechRole-Data.
+Finally, we conduct a human–LLM agreement study to validate the reliability of the proposed automatic evaluation framework.
+5.1 Evaluated Systems
+We evaluate a total of seven baseline SRPAs, grouped into cascaded and end-to-end systems.
+5.1.0.1 Cascaded Systems.
+Cascaded SRPAs follow a core three-stage pipeline consisting of ASR, text-based reasoning, and TTS.
+Specifically, user speech is first transcribed using whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23].
+The transcribed text is then processed by a large language model for generating role-conditioned responses.
+We consider three open-source LLMs for this stage: Qwen3-8B [DBLP:journals/corr/abs-2505-09388], Llama-3.1-8B [DBLP:journals/corr/abs-2407-21783], and Mistral-7B-Instruct-v0.3 [DBLP:journals/corr/abs-2310-06825].
+Finally, generated responses are converted into speech using F5-TTS [DBLP:journals/corr/abs-2410-06885].
+The TTS component supports voice cloning based on reference audio, enabling the synthesized speech to match the target character’s timbre.
+In addition to open-source cascaded systems, we include a proprietary cascaded baseline built upon the Alibaba Cloud API [alibabacloud], which integrates Paraformer Realtime ASR v2, Qwen-Plus-Character for role-conditioned text generation, and CosyVoice-v2 for speech synthesis.
+This system also supports reference-based voice cloning.
+5.1.0.2 End-to-End Systems.
+End-to-end SRPAs directly generate speech responses from speech inputs, avoiding explicit intermediate text representations.
+Such systems are expected to reduce error accumulation and latency, and are closer to the ideal formulation of SRPAs.
+We evaluate three end-to-end models: the proprietary GPT-4o Audio [DBLP:journals/corr/abs-2410-21276], and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] and LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025].
+Notably, LLaMA-Omni only supports English and is therefore excluded from Chinese evaluations.
+To further investigate the impact of role-specific speech data, we fine-tune Qwen2.5-Omni-7B on SpeechRole-Data, resulting in a specialized model referred to as SpeechRole-Agent.
+We compare its performance against the base model to assess the benefits of targeted SRPA training.
+5.2 Overall Results and System Comparison
+Table 2 summarizes the overall performance of all SRPAs on SpeechRole-Eval in English and Chinese.
+Overall, closed-source systems achieve the strongest results, with GPT-4o Audio consistently ranking at or near the top across languages.
+The cascaded system based on Alibaba Cloud API also performs competitively, particularly in Chinese.
+In contrast, open-source end-to-end models lag significantly behind across most metrics, resulting in substantially lower overall scores.
+5.2.0.1 Interaction Ability.
+IA and CC are strongly related to a model’s language understanding and reasoning ability, but our results show that they are also influenced by the speech modeling paradigm.
+Cascaded systems built on strong text-based LLMs, such as Qwen3-8B and Llama-3.1-8B, achieve consistently high scores on both metrics.
+However, end-to-end models may exhibit substantial degradation even when sharing the same text backbone.
+For example, LLaMA-Omni underperforms its cascaded Llama-3.1-8B counterpart on IA and CC, indicating that unified speech-text modeling can impair interaction quality if cross-modal integration is not sufficiently optimized.
+In contrast, GPT-4o Audio demonstrates that strong end-to-end SRPAs can still maintain robust interaction ability.
+Overall, these results suggest that IA and CC depend jointly on language modeling strength and effective speech-text integration.
+5.2.0.2 Speech Quality and Expressiveness.
+GPT-4o Audio demonstrates clear advantages in SF and SN, reflecting the benefits of tightly integrated speech generation.
+However, its gains in PC and EA are more limited, with several cascaded systems achieving comparable or even slightly stronger performance.
+This suggests that while end-to-end modeling improves basic perceptual speech quality, expressive control remains challenging.
+Overall, despite their theoretical advantages, current open-source end-to-end systems do not yet consistently surpass well-engineered cascaded approaches on expressive speech metrics.
+5.2.0.3 Role-Playing Fidelity.
+PeC and KC exhibit trends broadly aligned with interaction ability, with systems based on stronger language models generally maintaining more stable persona traits and role-consistent knowledge.
+However, similar to IA and CC, end-to-end models can show notable degradation even when built on comparable text backbones.
+For instance, LLaMA-Omni underperforms its cascaded counterpart in both PeC and KC, suggesting that unified speech-text modeling may weaken persona grounding and knowledge consistency when cross-modal alignment is insufficient.
+At the same time, GPT-4o Audio achieves competitive role fidelity, indicating that end-to-end designs can preserve character consistency when adequately optimized.
+Overall, role-playing fidelity depends on both the strength of the language model and the effectiveness of speech-text integration.
+5.3 Effect of Role-Specific Training Data
+To assess the impact of role-specific speech training, we compare the base Qwen2.5-Omni model with SpeechRole-Agent, which is fine-tuned on SpeechRole-Data covering 78 training roles.
+The evaluation is further divided into in-domain roles seen during training and out-of-domain roles that are held out, allowing us to examine both fitting and generalization behavior.
+The quantitative results are summarized in Table 3.
+SpeechRole-Agent achieves substantial improvements over the base model in the in-domain setting across all metrics, with particularly large gains in IA, CC, PeC, and KC.
+This indicates that role-specific supervision effectively enhances structured role-playing behavior in end-to-end systems.
+Importantly, these improvements largely persist in the out-of-domain setting.
+Although some degradation is observed compared to in-domain roles, SpeechRole-Agent consistently outperforms Qwen2.5-Omni across both languages and most evaluation dimensions, leading to significantly higher overall scores.
+This suggests that training on SpeechRole-Data improves not only role memorization but also general role-playing competence.
+5.4 Human–LLM Agreement Analysis
+To assess whether the proposed LLM-based evaluation aligns with human judgment, we conduct a human–LLM agreement study on a subset of SpeechRole-Eval.
+We focus on Chinese role-playing dialogues to facilitate consistent expert assessment under a shared linguistic context.
+Half of the Chinese evaluation instances are uniformly sampled, covering all dialogue types and role categories, and all baseline systems are evaluated on each instance.
+Three expert annotators independently assign absolute scores on a 1–10 scale following the evaluation dimensions defined in Section 4.1.
+Annotators are blinded to system identities and LLM scores.
+Human ratings are aggregated by averaging scores across annotators and evaluation instances for each baseline system.
+We evaluate agreement by computing Spearman’s rank correlation between the aggregated human scores and the automatic scores produced by SpeechRole-Eval at the system level.
+Strong correlations are observed across evaluation dimensions, indicating that the proposed LLM-based evaluation produces rankings aligned with human judgment and can serve as a proxy for expert evaluation.
+Detailed annotation protocols and additional analysis are provided in Appendix F.
+6 Conclusion
+This work introduces SpeechRole, a unified framework for advancing SRPAs through both large-scale data and systematic evaluation.
+We present SpeechRole-Data, a speech-to-speech role-playing dataset covering 98 roles with rich timbral and prosodic variation, and SpeechRole-Eval, a multidimensional benchmark that directly evaluates generated speech across interaction ability, speech expressiveness, and role-playing fidelity.
+Through extensive experiments on both cascaded and end-to-end systems, we provide a comprehensive analysis of current SRPAs.
+Our results show that while state-of-the-art end-to-end models achieve strong fluency and naturalness, challenges remain in prosody control and emotional expressiveness.
+Moreover, interaction ability and role-playing fidelity are largely constrained by the underlying text-based language models, highlighting a key bottleneck for future progress.
+We further demonstrate that the proposed LLM-based evaluation framework aligns well with expert human judgments, supporting its use as a scalable and reliable alternative to human evaluation.
+We hope SpeechRole will serve as a foundation for future research on expressive, controllable, and character-consistent speech role-playing systems.
+7 Limitations
+Despite its contributions, this work has several limitations.
+First, although SpeechRole-Data covers a diverse set of roles and speaking styles, the number of languages remains limited, with experiments primarily focusing on English and Chinese.
+Extending the dataset and evaluation framework to more languages and cultural contexts is an important direction for future work.
+Second, our automatic evaluation relies on a strong proprietary LLM as the judge.
+While human correlation analysis shows high agreement, the evaluation may still inherit biases or blind spots from the underlying model.
+Developing fully open-source judges or complementary objective speech metrics remains an open challenge.
+Third, the current evaluation focuses on system-level performance and averaged behavior.
+It does not explicitly assess fine-grained controllability, long-term character drift, or robustness under adversarial or ambiguous role instructions.
+These aspects are critical for real-world deployment of SRPAs and warrant further investigation.
+Finally, while SpeechRole-Agent demonstrates clear gains through task-specific training, it is built upon an existing end-to-end model.
+Exploring alternative architectures, training objectives, and tighter integration between language and speech modeling may further improve expressive and faithful role-playing in future systems.
+8 Ethics Statement
+This work does not involve the release of newly annotated human-labeled datasets.
+Human annotation is only conducted for evaluation purposes in the Human–LLM agreement study, and the resulting annotations are not included in the released resources.
+All human evaluations are performed by expert annotators selected from the research team, following a predefined evaluation protocol.
+No personal data is collected, and all evaluated content is derived from fictional role-playing dialogues.
+Therefore, we believe this work raises no significant ethical concerns.
+Appendix A AI Assistants in Research or Writing
+In preparing this manuscript, AI assistants were employed solely to assist with refining the clarity, style, and readability of certain text segments.
+They were not involved in designing the study, developing or implementing the methodology, collecting or analyzing data, or generating the primary scientific contributions.
+All substantive research decisions, analyses, and conclusions are fully the responsibility of the authors.
+Appendix B Evaluation Metrics Details
   • Instruction Adherence (IA): Do the spoken responses strictly follow the task instruction, remaining fully in character without any out-of-role explanations or assistant-like meta-comments?
+  • Conversational Coherence (CC): Do the responses maintain logical consistency within the dialogue, aligning with previous content without contradictions or abrupt topic shifts?
   • Speech Fluency (SF): Are the responses delivered fluently, with smooth articulation, appropriate pacing, and minimal disfluencies such as stuttering or unnatural pauses?
-  • Conversational Coherence (CC): Do the responses maintain logical consistency within the dialogue, aligning with previous content without contradictions or abrupt topic shifts?
-Speech quality and expressiveness measure the naturalness and prosodic expressiveness of the generated speech, which are assessed by the following metrics:
   • Speech Naturalness (SN): Do the responses sound natural, human-like, and free from noticeable artifacts or robotic effects typically associated with synthetic speech?
@@
   • Emotion Appropriateness (EA): Are emotional cues in the speech (e.g., anger, joy, sadness) well-aligned with the dialogue context and the character’s emotional state?
-Role-playing fidelity assesses the extent to which the agent embodies the assigned character’s persona, which is assessed by the following metrics:
   • Personality Consistency (PeC): Do the responses consistently reflect the character’s personality traits, such as optimism, sarcasm, or authority?
   • Knowledge Consistency (KC): Are the responses grounded in the character’s established background, knowledge, and relationships, without fabricating out-of-character facts?
-4.3 Evaluation Procedure
-LLM-Based Scoring Procedure.
-There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [DBLP:conf/acl/WangLCCZLCKLLS24].
-Inspired by DBLP:conf/iclr/DaiH0J0025, two strategies are adopted to enhance the reliability of LLM-based scoring.
-First, the model is required to generate a brief rationale for its score prior to providing the original rating.
-This process encourages more deliberate and interpretable evaluations by the model.
-Second, rather than evaluating individual speech samples in isolation, a paired comparison setting is designed.
-For each evaluation instance, the model receives a test speech sample and a high-quality reference answer simultaneously, and assigns scores to both.
-The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies.
-This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments.
-We use gemini-2.5-pro-preview-06-05 [DBLP:journals/corr/abs-2312-11805] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
-Human Correlation Verification.
-To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses.
-Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric.
-The final human score for each metric is calculated as the average of all annotators’ judgments.
-The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation.
-The human judgment setting follows the evaluation protocol introduced by DBLP:conf/iclr/DaiH0J0025, which uses pairwise preference comparisons to ensure stable and interpretable annotations.
-Details are in the appendix.
-Models Evaluation Metrics Overall
-IA SF $C$ SN PC EA PeC KC
-English Evaluation Results
-Alibaba Cloud API 0.9012 0.6678 1.0320 0.6973 0.6651 0.8087 0.8594 0.9171 0.8186
-XTTS 0.9617 0.5493 0.9872 0.6578 0.7568 1.0133 1.0357 0.9531 0.8644
-CosyVoice2 1.0102 0.7869 1.0346 0.8682 0.9857 1.0781 1.0476 0.9744 0.9732
-E2 TTS 1.0531 0.7938 1.1621 0.8859 1.0718 1.1806 1.1760 1.0104 1.0417
-F5-TTS 1.1245 0.8057 1.1132 0.8717 0.9773 1.2686 1.0952 1.0614 1.0397
-FireRedTTS 0.9587 0.7266 1.0297 0.8135 0.8760 1.0557 1.0415 0.9899 0.9365
-Spark-TTS 1.0384 0.7583 1.0641 0.8282 1.0554 1.1427 1.0741 1.0144 0.9970
-GPT-SoVITS 0.8566 0.4783 0.9293 0.5576 0.5606 0.8253 0.8899 0.8770 0.7468
-MaskGCT 0.9277 0.7635 0.9902 0.7589 0.9449 1.0382 0.9701 1.0098 0.9254
-[…]
-[…]
-Chinese Evaluation Results
-Alibaba Cloud API 0.8988 0.8473 0.8942 0.9275 1.0223 1.0493 0.9124 0.8887 0.9301
-XTTS 0.6563 0.4951 0.7149 0.5348 0.4681 0.4945 0.6191 0.6595 0.5803
-CosyVoice2 0.7600 0.7739 0.7190 0.7943 0.7388 0.8843 0.6587 0.6597 0.7486
-E2 TTS 0.6866 0.6513 0.7171 0.7082 0.6776 0.7027 0.6787 0.7161 0.6923
-F5-TTS 0.7285 0.6901 0.7658 0.8052 0.8392 0.9291 0.6969 0.6890 0.7680
-FireRedTTS 0.7536 0.6932 0.6968 0.7221 0.8371 0.8235 0.7093 0.6516 0.7359
-Spark-TTS 0.7591 0.7179 0.7869 0.7887 0.8930 0.7902 0.7268 0.6890 0.7690
-GPT-SoVITS 0.6915 0.6345 0.7269 0.6119 0.6583 0.6515 0.6296 0.6277 0.6540
-MaskGCT 0.7391 0.5913 0.6558 0.7349 0.8011 0.8109 0.7722 0.6433 0.7186
-[…]
-[…]
-Table 2: Comprehensive evaluation results of $1$ SRPAs on the SpeechRole-Eval benchmark.
-The table reports model performance across eight evaluation metrics: IA (Instruction Adherence), SF (Speech Fluency), $C$ (Conversational Coherence), SN (Speech Naturalness), PC (Prosody Consistency), EA (Emotion Appropriateness), PeC (Personality Consistency), and KC (Knowledge Consistency), along with the overall average score.
-Results are presented for both English and Chinese test sets.
-End-to-end models are indicated with a gray background.
-Proprietary models include Alibaba Cloud API and GPT-4o Audio, while all others are open-source.
-5 Experiments
-This section details our experimental setup and main results.
-We evaluate SRPAs under two paradigms: cascaded systems, which separate role-playing into ASR, LLM-based reasoning, and TTS stages, and end-to-end models that generate speech directly from prompts.
-All models are tested on the SpeechRole-Eval benchmark (Section 4) covering eight core metrics, with speaker similarity assessed for voice-controllable models.
-We also include supplementary experiments, such as fine-tuning a SRPA on our dataset to assess its speech role-playing performance.
-5.1 SRPA Architectures
-The primary objective of SRPAs is to generate coherent, spoken responses that are consistent with the target character’s identity upon receiving a user’s spoken query.
-To explore this, we examine two representative system architectures: the cascaded and end-to-end approaches.
-Cascaded SRPAs decompose the overall task into three modular components.
-First, an ASR model transcribes the user’s spoken input into text.
-Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role.
-Finally, a TTS module converts the generated text into speech that matches the target character’s voice.
-In our implementation, both proprietary and open-source cascaded systems are explored.
-We utilize a proprietary cascaded system based on the Alibaba Cloud API [alibabacloud], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2.
-For the open-source side, we employ whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23] for high-quality ASR, LLaMA3.1-8B [DBLP:journals/corr/abs-2407-21783] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], E2 TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], F5-TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], GPT-SoVITS [gpt-sovits], Spark-TTS [DBLP:journals/corr/abs-2503-01710], and MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025].
-End-to-end SRPAs generate speech responses directly from spoken input, guided by the user’s spoken query and a textual prompt of the character’s identity.
-Both the proprietary gpt-4o-audio-preview-2024-12-17 [DBLP:journals/corr/abs-2410-21276] and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] are evaluated in our study.
-5.2 Evaluation Results
-Main Evaluation Results.
-We report the main evaluation results of all SRPAs on the SpeechRole-Eval benchmark.
-This benchmark covers eight core metrics: Instruction Adherence (IA), Speech Fluency (SF), Conversational Coherence (CC), Speech Naturalness (SN), Prosody Consistency (PC), Emotion Appropriateness (EA), Personality Consistency (PeC), and Knowledge Consistency (KC), with the final score computed as the average across these dimensions.
-As shown in Table 2, the cascaded agents generally outperform end-to-end systems across most evaluation dimensions.
-Notably, models such as E2 TTS, F5-TTS, and Spark-TTS achieve relatively high overall scores in both content relevance and expressiveness, benefiting greatly from their modular design and optimization.
-While end-to-end models, particularly GPT-4o Audio, demonstrate competitive performance in expressiveness and coherence, they still exhibit noticeable shortcomings in style consistency and character setting.
-The performance of Qwen2.5-Omni is significantly lower across all metrics, indicating that end-to-end modeling remains challenging in the absence of sufficient task-specific supervision.
-Language-Level Observations.
-We observe significant differences between the English and Chinese subsets.
-Overall, the model performs better on English, particularly with respect to metrics such as content coherence and knowledge consistency.
-This disparity likely reflects the model’s language imbalance in pre-training and TTS capabilities.
-For example, XTTS performs relatively well in English but shows a marked decline in Chinese, especially in the expression of emotion and character personality.
-In contrast, Alibaba Cloud API and CosyVoice2 demonstrate more stable performance across both languages.
-These results indicate that current SRPAs exhibit uneven performance across different languages, with significantly better results in English than in Chinese.
-Therefore, further efforts are necessary to develop high-quality multilingual SRPAs.
-Figure 4: Speaker similarity scores of SRPAs with voice cloning capability.
-Speaker Similarity Comparison.
-We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity.
-Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems.
-Following the same evaluation protocol as Seed-TTS [DBLP:journals/corr/abs-2406-02430], we use the WavLM-large [DBLP:journals/jstsp/ChenWCWLCLKYXWZ22] model to extract speaker embeddings and compute cosine similarity scores.
-As shown in Figure 4, in the English setting, E2 TTS and F5-TTS achieve the highest speaker similarity, while GPT-SoVITS performs the worst.
-The overall ranking of models remains largely consistent across the two languages, indicating stable voice cloning performance.
-Notably, MaskGCT slightly outperforms others in Chinese, reaching the highest SIM score.
-Models Evaluation Metrics Overall
-IA SF $C$ SN PC EA PeC KC
-English Evaluation Results
-Qwen2.5-Omni 0.3977 0.4968 0.5491 0.4505 0.2958 0.3814 0.4089 0.4917 0.4340
-SpeechRole-Agent 0.8072 0.6790 0.9480 0.6123 0.4202 0.5383 0.7568 0.9169 0.7098
-[…]
-[…]
-Chinese Evaluation Results
-Qwen2.5-Omni 0.4924 0.6886 0.5832 0.6204 0.4693 0.5125 0.3932 0.4772 0.5296
-SpeechRole-Agent 0.9188 0.8233 0.9037 0.7749 0.7277 0.8607 0.8402 0.9253 0.8468
-[…]
-[…]
-Table 3: Performance comparison between Qwen2.5-Omni and the fine-tuned SpeechRole-Agent on the SpeechRole-Eval benchmark.
-Results are reported for both English and Chinese, covering in-domain and out-of-domain scenarios.
-Out-of-domain results are indicated with gray background.
-5.3 Fine-Tuning SRPA
-Training Setup.
-We fine-tune Qwen2.5-Omni-7B on the SpeechRole training set, containing 78 roles with speech-based dialogues.
-The model trains using speech inputs paired with role-consistent responses, enabling it to better understand and generate role-based spoken interactions.
-During evaluation, we use SpeechRole-Eval, dividing it into two subsets: in-domain, which includes roles seen during training, and out-of-domain, consisting of unseen roles.
-Results are reported separately for English and Chinese.
-Performance Analysis.
-After fine-tuning on the complete SpeechRole-Data training set, SpeechRole-Agent achieves notable gains over the original Qwen2.5-Omni.
-In the in-domain test set, the overall score in English increases from 0.4340 to 0.7098, while in Chinese it rises from 0.5296 to 0.8468.
-All evaluation metrics improve, with particularly notable advancements in instruction adherence, emotional appropriateness, and personality consistency.
-In the out-of-domain test set, the model also demonstrates strong generalization ability, with the English score increasing from 0.4786 to 0.6795 and the Chinese score from 0.5117 to 0.8028.
-These results fully demonstrate the model’s robust role-playing capabilities beyond the scope of the training set.
-Detailed results are presented in Table 3.
-6 Conclusion
-We propose SpeechRole, a unified framework for developing and evaluating SRPAs.
-To support this objective, we construct SpeechRole-Data, a large-scale, high-quality dataset comprising 112k dialogues from 98 characters, each with rich role settings and diverse vocal characteristics.
-Our benchmark establishes rigorous evaluation protocols for role-consistent speech generation and introduces SpeechRole-Eval, an automated multimodal evaluation suite that assesses various aspects including interaction competence, speech expressiveness, and role consistency.
-Experimental results indicate current SRPAs still face challenges in achieving fully coherent and expressive role simulation, underscoring the need for more advanced character-based speech behavior modeling.
-We expect this work will provide a solid foundation for future research at the intersection of speech generation, role-playing, and embodied agents.
-All data, evaluation protocols, and baseline methods will be made publicly available to facilitate further studies on speech-interactive agents.
-Figure 5: Left: Distribution of eight sub-tasks across three major task categories in SpeechRole. Right: Distribution of the 98 roles in SpeechRole-Data by language, gender, and data source.
-Figure 6: Main evaluation results of $1$ SRPAs on SpeechRole-Eval.
-Left: English results.
-Right: Chinese results.Each metric is displayed with an interval range of 0.26, with all evaluation results normalized such that the maximum value across metrics is scaled to 1.30.
-Appendix A SpeechRole-Data Construction
-A.1 Text Data Construction
-Role Selection.
-A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024].
-Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM.
-The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5).
-To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics.
-Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
-Role Metadata Extraction.
-For each character, we extract structured metadata, including: (1) Role Profile, such as temperament, behavioral tendencies, values, and personal preferences; (2) Background, including character settings, social identity, relationships, and living environment; and (3) Character Lines, including narration, monologues, and dialogue excerpts involving the target character in games or audiovisual works.
-These elements collectively define distinctive character profiles and support vivid, context-rich speech role-playing.
-Additionally, the dialogue excerpts serve as the basis for guiding dialogue generation and evaluating character consistency.
-Dialogue Generation.
-We use gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate dialogues between users and roles based on role metadata.
-For some roles with lengthy scripts, the content is divided into segments and processed sequentially.
-GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions.
-For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
-Dialogue Filtering.
-The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity.
-First, only dialogues that conform to a clear turn-taking pattern are retained.
-In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns.
-To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts.
-For English data, the all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] model is used, and samples with a similarity score below 0.4 are excluded.
-The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [text2vec] model, with a similarity score threshold of 0.45.
-To further eliminate semantic redundancy, which is commonly observed in GPT-4.1-generated dialogues, near-duplicate samples are removed based on the similarity of their dialogue embeddings.
-Thresholds are set to 0.85 for Chinese and 0.9 for English.
-This de-duplication step is conducted separately for single-turn and multi-turn dialogues for each role.
-After filtering, the dataset contains 53,902 single-turn dialogues and 57,944 multi-turn dialogues.
-Metrics IA SF $C$ SN PC EA PeC KC Overall
-MSE 0.1020 0.0413 0.1426 0.0581 0.2539 0.1262 0.1690 0.1147 0.1260
-Pearson 0.7304 0.6366 0.6076 0.8093 0.6316 0.8881 0.6677 0.7581 0.7162
-Table 4: Evaluation results of correlation between automatic and human scoring.
-A.2 Role Voice Collection
-Role Voice Extraction.
-To achieve authentic voice role-playing, real-life performance audio for each character is collected as reference material.
-For characters from Genshin Impact, in-game character audio is collected with permission from miHoYo.
-For other characters, audio is obtained from actual film and television works.
-Films and television series featuring these characters are gathered, and a total of 105 video files with an approximate duration of 181 hours are downloaded from the internet, covering all target characters.
-During the preprocessing stage, all extracted speech segments undergo anonymization and fragmentation to ensure that they contain only brief, decontextualized utterances, with no retention of the original video or narrative context.
-As a result, no identifiable or copyrighted content remains, thus mitigating potential intellectual property risks and ensuring adherence to the principles of fair use for research purposes.
-We first employ a batch processing pipeline based on ffmpeg [ffmpeg] to extract the raw audio streams from the collected video files.
-To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format.
-However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
-Role Voice Processing.
-To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [DBLP:conf/slt/HeSWLGHLYLSWCZW24] framework.
-This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation.
-The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
-Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22] metric.
-In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3.
-The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
-Speaker Identification.
-The extracted audio data contains only the speaker’s index and does not include identity information.
-To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437] models to determine the target character’s speaker number based on all transcribed texts.
-We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
-Role Voice Ranking.
-To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity.
-Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [DBLP:conf/interspeech/WangZCC023] model and calculate the cosine similarity between each sample and all other samples.
-A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role.
-Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.
-Appendix B Human Correlation Verification Details
-To evaluate the consistency between automatic scoring and human scoring, we randomly selected 20 samples from the dataset, including 5 single-turn dialogue samples and 15 multi-turn dialogue samples.
-For each sample, we randomly selected two model-generated speech responses for evaluation.
-Gemini scored each model response and its corresponding reference response on a scale of 1 to 10 based on 8 predefined scoring metrics.
-The final score for a model was calculated as the ratio of the model response score to the reference response score.
-For human evaluation, four trained evaluators were employed to directly compare the responses of the two models based on the same 8 scoring metrics.
-The evaluation criteria were as follows: the superior model received 1 point, the inferior model 0 points, and if the performance was deemed equivalent, each received 0.5 points.
-The final human score was determined by averaging the scores from all evaluators.
-To quantify the agreement between automated and human evaluation methods, we computed two key metrics: mean squared error (MSE) and Pearson correlation coefficient.
-The MSE was calculated by first fitting a linear regression between the score differences of paired model responses from human evaluators and gemini.
-The predicted gemini score differences were then derived from this regression model, and the MSE was computed between these predicted values and the actual gemini score differences.
-For the Pearson correlation analysis, we computed the linear correlation coefficient between the human evaluators’ score differences and gemini’s score differences across all paired comparisons.
-Table 4 reports the detailed results, the overall MSE of 0.1260 indicates relatively minor deviations between the two scoring methods.
-The high Pearson correlation coefficient further confirms the reliability of the automated scoring system.
-The evaluation results collectively demonstrate strong agreement between automatic and human scoring across all metrics, indicating that the automated scoring system can effectively approximate human judgment.
 Appendix C Additional data statistics
-Role Type Distribution
+C.0.0.1 Role Type Distribution
 Figure 5 presents the distribution of characters by language (Chinese and English), source type (game, film, television series, and animation), and gender.
@@
 This diversity provides extensive coverage of voices and contexts for role-playing.
-Task Type Distribution
+C.0.0.2 Task Type Distribution
 User prompts can generally be categorized into three major task categories, which can be further divided into eight subtask types, as illustrated in the figure 4.
@@
     • Advice and Life Experience: Questions seeking advice, coping methods, or life experiences. (13.61%)
-Evaluation Results Demonstration
-Figure 6 visually presents the evaluation results of all SRPAs on the SpeechRole-Eval benchmark.
+Table 4: Average Human Evaluation Scores across Models
+Metric Alibaba Cloud API GPT-4o Audio Llama-3.1-8B Mistral-7B Qwen2.5-Omni Qwen3-8B
+Instruction Adherence (IA) 8.61 7.99 7.79 7.71 7.72 7.30
+Conversational Coherence (CC) 8.52 7.54 7.17 7.07 7.19 6.58
+Speech Fluency (SF) 8.50 7.06 6.47 6.78 7.26 6.04
+Speech Naturalness (SN) 8.40 6.75 7.08 6.99 6.67 6.46
+Prosody Consistency (PC) 8.49 6.69 7.60 7.34 3.70 7.16
+Emotion Appropriateness (EA) 8.41 5.33 7.49 7.29 4.81 6.96
+Personality Consistency (PeC) 8.46 5.84 7.59 7.09 5.18 7.08
+Knowledge Consistency (KC) 8.25 7.40 7.48 6.89 6.97 7.19
+Table 5: Spearman correlation between human judgments and automatic evaluation.
+Metric Spearman $\rho$
+Speech Fluency (SF) 0.771
+Instruction Adherence (IA) 0.600
+Conversational Coherence (CC) 0.543
+Knowledge Consistency (KC) 0.486
+Speech Naturalness (SN) 0.371
+Prosody Consistency (PC) 0.371
+Emotion Appropriateness (EA) 0.371
+Personality Consistency (PeC) 0.371
 Appendix D Prompt Templates
-Prompts for Batch Data Generation
+D.0.0.1 Prompts for Batch Data Generation
 Figure 6 and Figure 7 respectively present the prompts used for generating single-turn dialogue data for English and Chinese characters.
 Figure 8 and Figure 9 respectively present the prompts used for generating multi-turn dialogue data for English and Chinese characters.
-Prompts for Automated Judgement
+D.0.0.2 Prompts for Automated Judgement
 Figure 10 displays the judgment prompt designed for automatically assessing audio responses generated by the two models.
 Appendix E SpeechRole-Agent training settings
-Figure 12 presents the distributed training script for fine-tuning the Qwen2.5-Omni-7B model on 8×H20 GPUs.
+Figure 11 presents the distributed training script for fine-tuning the Qwen2.5-Omni-7B model on 8×H100 GPUs.
 The configuration employs bfloat16 mixed-precision training with a learning rate of 1e-4 and gradient accumulation.
 The implementation includes periodic evaluation (every 500 steps) and checkpoint management.
+Appendix F Human–LLM Agreement Analysis Details
+We conduct a human–automatic alignment analysis on the Chinese subset of SpeechRole-Eval.
+Three internal expert annotators independently evaluate single-turn and multi-turn dialogues generated by six models across sixteen roles, following unified written instructions.
+All annotators are internal researchers and participate voluntarily without monetary compensation.
+The evaluated dialogues are fictional role-playing content used solely for research and benchmarking purposes.
+Human scores are aggregated at the system level and compared with Gemini-based automatic rankings using Spearman’s rank correlation.
+As shown in Table 4 and Table 5, positive correlations are observed across all eight evaluation metrics, with Spearman’s $\rho$ exceeding 0.37 for every metric, indicating that the proposed automatic evaluation produces system-level rankings broadly consistent with expert human judgments.
 Role Name Category Language Gender Source Split